x-12/7+5/x/1/2-7/2x
can you simplify this
5/x+7/9=-5/x
can you determine whether this is a equation or expression and solve it
2x/x+7-1/x+1=-2x+8/x^2-8x+7
can you solve this equation

Answer by jcmtnez(53)


(1) 5/x+7/9=-5/x, This is equal to 45+7x=-45, 7x=50, x=50/7
(2) 2x/x+7-1/x+1=-2x+8/x^2-8x+7 I have change the sign of 8x to make it possible factorization.
2x/x+7-1/x+1=-2x+8/x^2-8x+7
2x/x+7-1/x+1=-2x+8/((x+7)(x+1))
2x(x+1)-(x+7)=2x+8
2x^2+2x-x-7=2x+8
2x^2-x-15=0
Know using the formula we get the solutions: x=10/4 and x=-3
Remember that I have change on sign.